### v2

# GIST: Greedy Independent Set Thresholding for Diverse Data Summarization

Abstract
We introduce a novel subset selection problem called min-distance diversification with monotone submodular utility (MDMS), which has a wide variety of applications in machine learning, e.g., data sampling and feature selection. Given a set of points in a metric space, the goal of MDMS is to maximize an objective function combining a monotone submodular utility term and a min-distance diversity term between any pair of selected points, subject to a cardinality constraint. We propose the GIST algorithm, which achieves a $\nicefrac{{1}}{{2}}$-approximation guarantee for MDMS by approximating a series of maximum independent set problems with a bicriteria greedy algorithm. We also prove that it is NP-hard to approximate to within a factor of $0.5584$. Finally, we demonstrate that GIST outperforms existing benchmarks for on a real-world image classification task that studies single-shot subset selection for ImageNet.
1 Introduction
Subset selection is a challenging combinatorial optimization problem with a wide variety of applications in machine learning, including feature selection, recommender systems, drug discovery, data summarization, and designing pretraining sets for large language models (Team et al., 2024). Data sampling in particular is a salient problem due to unprecedented and continuous streams of data collection. For example, LiDAR and imaging devices in one self-driving vehicle can easily capture ~80 terabytes of data per day (Kazhamiaka et al., 2021).
In many subset selection tasks, we rely on the utility (or weight) of an object to rank one over the other, and also to avoid selecting duplicate or near-duplicate objects. If we select a small subset, then we must ensure that the selected set is a good representation of the original set. These utility, diversity, and coverage criteria are expressed through objective functions, and the interesting research lies in developing efficient algorithms with strong approximation guarantees. The algorithmic tools used in constrained subset selection shares many similarities with techniques from other areas of combinatorial optimization such as submodular maximization, $k$-center clustering, and convex hull approximations.
This work introduces a novel subset selection problem called min-distance diversification with monotone submodular utility (MDMS). Given $n$ points $V$ in a metric space, a monotone submodular function $g:2^{V}\rightarrow\mathbb{R}_{\geq 0}$, and a cardinality constraint $k$, the goal is to solve
$S^{*}$ $=\operatorname*{arg\,max}_{S\subseteq V}~{}g(S)+\texttt{div}(S)$ (1)
subject $\hskip 13.6572pt\text{subject to}~{}~{}|S|\leq k,$
where the diversity term $\texttt{div}(S)=\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$ encourages maximizing the min distance between any two distinct points in the selected set. Meinl et al. (2011) call $\texttt{div}(S)$ the $p$-dispersion measure, and Bhaskara et al. (2016) call it the min-min diversity objective, where is known to admit a $\nicefrac{{1}}{{2}}$-approximation ratio subject to the strict cardinality constraint $|S|=k$.
Other diversity terms such as $\sum_{u\in S}\min_{v\in S\setminus\{u\}}\textnormal{dist}(u,v)$ and $\sum_{u\in S}\sum_{v\in S}\textnormal{dist}(u,v)$ have also been studied (Meinl et al., 2011; Bhaskara et al., 2016; Borodin et al., 2017). While previous works have explored maximizing $\texttt{div}(S)$ in isolation, MDMS introduces a utility term and has not been studied before. Bhaskara et al. (2016) remark that $\texttt{div}(S)$ is a powerful diversity objective, but point out a major challenge: it is highly non-monotonic, i.e., it can exhibit a stark decrease in the objective value by adding a single new point. Our work shows combining a monotone submodular utility term with $\texttt{div}(S)$ mitigates some of the challenges associated with this diversity measure, making it easier to leverage its properties.
1.1 Our Contributions
We summarize our main contributions below:
  • In Section 2, we formalize the MDMS problem and present a simple $0.387$-approximation algorithm as a warm-up, demonstrating how the competing terms in the objective function interact.
  • In Section 3, we present the GIST algorithm, which achieves a $\nicefrac{{1}}{{2}}$-approximation guarantee for MDMS by approximating a series of maximum independent set problems with a bicriteria greedy algorithm and returning the best solution. In the special case of linear utility functions $g(S)=\sum_{v\in S}w(v)$, we prove that GIST offers a stronger $\nicefrac{{2}}{{3}}$-approximation guarantee.
[…]
  • In Section 5, our experiments show that GIST outperforms baselines for MDMS on synthetic data (in particular the classic greedy algorithm). Then we show that GIST can be used to build better single-shot subsets of training data for an image classification benchmark on ImageNet compared to margin sampling and $k$-center algorithms, demonstrating the benefit of optimizing for a blend of utility and diversity.
1.2 Related Work
Diversity objectives.
A closely related problem is $k$-center, which aims to find a subset of $k$ centers that minimize the largest distance of any point to its nearest center in $S$. Concretely, the objective is $\max_{v\in V}\min_{c\in S}\textnormal{dist}(v,c)$, and a greedy $2$-approximation algorithm is widely used in active learning applications (Sener and Savarese, 2018). The main difference between MDMS and $k$-center is the ability to benefit from the utility term, e.g., uncertainty values in an active learning setting. The second key difference is that we relax the requirement to cover every point in the dataset, which can be helpful for large datasets where we want to avoid instability due to certain corner or outlier points.
Bhaskara et al. (2016) study various distance-based diversity metrics subject to a cardinality constraint. In particular, they give a $\nicefrac{{1}}{{8}}$-approximation algorithm for the sum-min diversity objective defined as $\sum_{u\in S}\min_{v\in S\setminus\{u\}}\textnormal{dist}(u,v)$ by rounding a linear programming relaxation. They complement this result by proving a conditional $\nicefrac{{1}}{{2}}$-hardness of approximation under the planted clique assumption. They also study the min-min objective $\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$, but they enforce a strict cardinality constraint $|S|=k$; otherwise, it is optimal to only output one or two points. Meinl et al. (2011) study the same objective subject to $|S|=k$ under the name $p$-dispersion (also called max-min or remote-edge) diversity maximization. This exact cardinality issue does not arise in MDMS because we need to balance the trade-off between competing monotone submodular utility and diversity terms.
Moumoulidou et al. (2020), Addanki et al. (2022), and Mahabadi and Trajanovski (2023) study diversity maximization under fairness constraints. Given a partitioning of the dataset, they ensure that a specified number of points $k_{i}$ are selected from each group $i\in[m]$. This fairness constraint considerably differentiates their problem setting from ours (i.e., $|S|=k=\sum_{i=1}^{m}k_{i}$). Moumoulidou et al. (2020); Addanki et al. (2022) maximize $\texttt{div}(S)$ in this setting. Mahabadi and Trajanovski (2023) consider sum-of-all pairwise distances $\sum_{u\in S}\sum_{v\in S}\textnormal{dist}(u,v)$ and sum-min objectives. However, none of these works add a utility term to the objective.
Borodin et al. (2017) optimize the combination of a submodular utility and sum-of-all pairwise distances in the selected set, making it the most similar work to MDMS. Unlike our $\texttt{div}(S)$ term, max-sum diversity is not very sensitive to duplicate points. In fact, adding $m$ copies of the same point to a set raises the increment of the diversity term by a factor of $m$ in their model. By using $\texttt{div}(S)$ instead, we can strictly avoid duplication across all selected points.
Data sampling.
In the realm of dataset curation and active learning, there have been several lines of work that study the combination of utility and diversity terms. Ash et al. (2020) invoke $k$-means++ seeding over the gradient space of a model in order to balance uncertainty and diversity. Wei et al. (2015) introduce several submodular objectives, e.g., facility location, and use them to diversify a set of uncertain examples in each active learning iteration. Citovsky et al. (2021) cluster examples that are represented by embeddings extracted from the penultimate layer of a partially trained DNN, and use these clusters to diversify uncertain examples in each iteration. Our work differs from these and several others (e.g., Kirsch et al. (2019); Zhdanov (2019)) in that we directly incorporate the utility and diversity terms in our objective function.
2 Preliminaries
Here we formalize the MDMS problem and present a simple $0.387$-approximation algorithm as a starting point.
Submodular set function.
For a set function $g:2^{V}\rightarrow\mathbb{R}$ and any $S,T\subseteq V$, let $g(T\mid S)=g(S\cup T)-g(T)$ denote the marginal gain of $g$ at $T$ with respect to $S$. A function $g$ is submodular if for every $S\subseteq T\subseteq V$ and $v\in V\setminus T$, we have $g(v\mid S)\geq g(v\mid T)$, where we overload the marginal gain notation for singleton sets. A submodular function $g$ is monotone if for every $S\subseteq T\subseteq V$, $g(S)\leq g(T)$.
[…]
Min-distance diversity.
For a set of $n$ points $V$ in a metric space, define the min-distance diversity function
$\texttt{div}(S)$ =$=\begin{cases}\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)&\text{if }|S|\geq 2,\\ \max_{u,v\in V}\textnormal{dist}(u,v)&\text{if }|S|\leq 1.\end{cases}$
For $\texttt{div}(S)$ to be well-defined, we set the diversity to be the diameter of $V$ if $|S|\leq 1$, i.e., $d_{\max}=\max_{u,v\in V}\textnormal{dist}(u,v)$. This implies that it is always beneficial to select at least two points to maximize the overall objective.
We extend the distance function to take subsets $S\subseteq V$ as input, i.e., $\textnormal{dist}(u,S)=\min_{v\in S}\textnormal{dist}(u,v)$. For the sake of completeness, we define $\textnormal{dist}(u,\varnothing)=\infty$.
Optimization problem.
Define the objective function
$f(S)=g(S)+\texttt{div}(S).$ (2)
The goal of the min-distance diversification with monotone submodular utility (MDMS) problem is to maximize $f(S)$ subject to a cardinality constraint $k$:
$S^{*}=\operatorname*{arg\,max}_{S\subseteq V:|S|\leq k}f(S).$
Let $\textnormal{OPT}=f(S^{*})$ denote the optimal objective value.
Remark 2.1.
The objective function $f(S)$ is not submodular. We present a counterexample in Appendix A.
Intersection graph.
Let $G_{d}(V)$ be the intersection graph of $V$ for distance threshold $d$, i.e., with nodes $V$ and edges
$E=\{(u,v)\in V^{2}:\textnormal{dist}(u,v)<d\}.$
Note that for any independent set $S$ of $G_{d}(V)$ and pair of distinct nodes $u,v\in S$, we have $\textnormal{dist}(u,v)\geq d$.
2.1 Warm-up: Simple $0.387$-Approximation Algorithm
The objective function $f(S)$ has two terms, each of which is easy to (approximately) optimize in isolation. For the $g(S)$ term, run the greedy algorithm to get $S_{1}$ satisfying
$g(S_{1})$ $\geq(1-\nicefrac{{1}}{{e}})\cdot\max_{|S|\leq k}g(S)$
$\geq(1-\nicefrac{{1}}{{e}})\cdot g(S^{*}).$
For the $\texttt{div}(S)$ term, let $S_{2}$ be a pair of diametrical points in $V$. The algorithm then flips a biased coin and outputs $S_{1}$ with probability $p\approx 0.613$, and $S_{2}$ otherwise. This achieves a $0.387$-approximation guarantee:
$\mathbb{E}[\textnormal{ALG}_{\textnormal{simple}}]$ $\geq p\cdot g(S_{1})+(1-p)\cdot\texttt{div}(S_{2})$
$\geq p\cdot(1-\nicefrac{{1}}{{e}})\cdot g(S^{*})+(1-p)\cdot\texttt{div}(S^{*})$
$=\frac{e-1}{2e-1}\cdot\textnormal{OPT},$
where we take $p=e/(2e-1)$ by solving $p\cdot(1-\nicefrac{{1}}{{e}})=1-p$. One of the main goals of this paper is to improve over this baseline and prove complementary hardness results.
3 Algorithm
In this section, we present the GIST algorithm and prove that it achieves an approximation ratio of $(\nicefrac{{1}}{{2}}-\varepsilon)$ for MDMS. GIST works by sweeping over multiple distance thresholds $d$ and calling a GreedyIndependentSet subroutine to find a high-value independent set of the intersection graph $G_{d}(V)$. This is in contrast to the simple $0.387$-approximation algorithm that just tries two extreme thresholds $D=\{0,d_{\max}\}$.
The GreedyIndependentSet heuristic builds a set $S$ (starting from the empty set) by repeatedly adding $v\in V\setminus S$ with the highest marginal gain with respect to the submodular function $g$ while satisfying $\textnormal{dist}(v,S)\geq d$. This subroutine runs until either $|S|=k$ or $S$ is a maximal independent set of $G_{d}(V)$.
[…]
Algorithm 1 Diverse data summarization for submodular utilities using greedy-weighted maximal independent sets.
1:function GIST(points $V$, monotone submodular function $g:V\rightarrow\mathbb{R}_{\geq 0}$, budget $k$, error $\varepsilon$)
2:     Initialize $S\leftarrow\textnormal{{GreedyIndependentSet}}(V,g,0,k)$triangleright runs the classic greedy algorithm
3:     Let $d_{\max}=\max_{u,v\in V}\textnormal{dist}(u,v)$ be the diameter of $V$
4:     Let $T\leftarrow\{u,v\}$ be two points such that $\textnormal{dist}(u,v)=d_{\max}$
5:     if $f(T)>f(S)$ and $k\geq 2$ then
6:         Update $S\leftarrow T$
7:     Let $D\leftarrow\{(1+\varepsilon)^{i}\cdot\varepsilon d_{\max}/2:(1+\varepsilon)^{i}\leq 2/\varepsilon\mbox{ and }i\in\mathbb{Z}_{\geq 0}\}$triangleright distance thresholds
8:     for threshold $d\in D$ do
9:         Set $T\leftarrow\textnormal{{GreedyIndependentSet}}(V,g,d,k)$
10:         if $f(T)\geq f(S)$ then
11:              Update $S\leftarrow T$               return $S$
1:function GreedyIndependentSet(points $V$, monotone submodular function $g:V\rightarrow\mathbb{R}_{\geq 0}$, distance $d$, budget $k$)
2:     Initialize $S\leftarrow\varnothing$
3:     for $i=1$ to $k$ do
4:         Let $C\leftarrow\{v\in V\setminus S:\textnormal{dist}(v,S)\geq d\}$
5:         if $C=\varnothing$ then
6:              return $S$triangleright $S$ is a maximal independent set of $G_{d}(V)$
7:         Find $t\leftarrow\operatorname*{arg\,max}_{v\in C}g(v\mid S)$
8:         Update $S\leftarrow S\cup\{t\}$      return $S$
Theorem 3.1.
For any $\varepsilon>0$, the GIST algorithm outputs a set $S\subseteq V$ with $|S|\leq k$ and $f(S)\geq(\nicefrac{{1}}{{2}}-\varepsilon)\cdot\textnormal{OPT}$.
The main building block in our design and analysis of GIST is the following lemma, which is inspired by the greedy $2$-approximation algorithm for metric $k$-centers and adapted for monotone submodular rewards.
Lemma 3.2.
Let $S_{d}^{*}$ be a maximum-value set of size at most $k$ with diversity at least $d$. In other words,
$S_{d}^{*}=\operatorname*{arg\,max}_{S:|S|\leq k,\texttt{div}(S)\geq d}g(S).$
[…]
Proof.
Let $k^{\prime}=|T|$ and $t_{1},t_{2},\dots,t_{k^{\prime}}$ be the points in $T$ in selection order. Let $B_{i}=\{v\in V:\textnormal{dist}(t_{i},v)<d^{\prime}\}$ be the points in $V$ in the radius-$d^{\prime}$ open ball around $t_{i}$. Since the distance between any pair of points in $B_{i}$ is at most $2d^{\prime}<d$, each set $B_{i}$ contains at most one point in $S_{d}^{*}$.
We first construct an injective map $h:S_{d}^{*}\rightarrow T$. We say that a point $v\in V$ is covered by $B_{i}$ if $v\in B_{i}$. For each covered $s\in S_{d}^{*}$, we map it to $t_{i}$, where $i$ is the minimum index for which $B_{i}$ covers $s$. If there is an $s\in S_{d}^{*}$ not covered by any set $B_{i}$, then GreedyIndependentSet must have selected $k$ points, i.e., $|T|=k$. We map the uncovered points in $S_{d}^{*}$ to arbitrary points in $T$ while preserving the injective property.
Since $g$ is a monotone submodular function, we have
$g(S_{d}^{*})-g(T)\leq\sum_{s\in S_{d}^{*}}g(s\mid T).$
We use $h$ to account for $g(s\mid T)$ in terms of the values we gained in set $T$. For any $s\in S_{d}^{*}$, let $T_{s}$ be the set of points added to set $T$ in algorithm GreedyIndependentSet right before the addition of $h(s)$. Since GreedyIndependentSet incrementally adds points and never deletes points from $T$, we know $T_{s}\subseteq T$. By submodularity, $g(s\mid T)\leq g(s\mid T_{s})$. By the greedy nature of the algorithm, we also know that
$g(h(s)\mid T_{s})\geq g(s\mid T_{s}).$
We note that the sum of the former terms, $g(h(s)\mid T_{s})$, is at most $g(T)$ since $h$ is injective and $g$ is nondecreasing. Thus, we conclude that
$g(S_{d}^{*})-g(T)$ $\leq\sum_{s\in S_{d}^{*}}g(s\mid T)$
$\leq\sum_{s\in S_{d}^{*}}g(h(s)\mid T_{s})$
$\leq g(T),$
which completes the proof. ∎
Equipped with this bicriteria approximation guarantee, we can now analyze the approximation ratio of GIST.
Proof of Theorem 3.1.
Let $d^{*}$ be the minimum pairwise distance between points in $S^{*}$. The GIST algorithm iterates over a set of thresholds $D$. The definition of $D$ implies that at least one threshold $d$ is in the range $[d^{*}/(2(1+\varepsilon)),d^{*}/2)$ unless $d^{*}\leq\varepsilon d_{\max}$. We deal with this special case later and focus on the case when such a $d$ exists.
Let $T$ be the output of GreedyIndependentSet for threshold value $d$. Applying Lemma 3.2 since $d<d^{*}/2$, we know that
$g(T)\geq\frac{1}{2}\cdot g(S_{d^{*}}^{*})\geq\frac{1}{2}\cdot g(S^{*}).$
We also know from our choice of $d$ that
$\texttt{div}(T)\geq\frac{1}{2(1+\varepsilon)}\cdot\texttt{div}(S^{*}).$
Combining these inequalities gives us
$f(T)\geq\frac{1}{2(1+\varepsilon)}\cdot\textnormal{OPT}>\left(\frac{1}{2}-\varepsilon\right)\cdot\textnormal{OPT}.$
It remains to prove the claim for the special case of $d^{*}\leq\varepsilon d_{\max}$. Algorithm GIST considers two initial feasible sets and picks the better of the two as the initial value for $T$. The first set is the classic greedy solution for the submodular function $g$, ignoring the diversity term. This gives the lower bound (Nemhauser et al., 1978):
$f(T)\geq\left(1-\frac{1}{e}\right)\cdot g(S^{*})\geq\left(1-\frac{1}{e}\right)\cdot\left(\textnormal{OPT}-\varepsilon d_{\max}\right).$
The other initial set realizes the max pairwise distance $d_{\max}$ and ignores the submodular term, yielding the lower bound:
$f(T)\geq d_{\max}.$
Combining these two lower bounds, we get a final bound of
$f(T)\geq\left(1-\frac{1}{e}\right)\cdot\frac{\textnormal{OPT}}{1+\varepsilon},$
which completes the proof. ∎
Remark 3.3.
We can modify GIST to guarantee an exact $\nicefrac{{1}}{{2}}$-approximation ratio by trying all pairwise thresholds $D^{\prime}=\{\textnormal{dist}(u,v)/2-\varepsilon^{\prime}:u,v\in V\}$, for a small enough $\varepsilon^{\prime}>0$, at the cost of $O(n^{2})$ calls to GreedyIndependentSet instead of $O(\log_{1+\varepsilon}(1/\varepsilon))$.
3.1 Linear Utility
We now assume $g(S)=\sum_{v\in S}w(v)$ is a linear function for some weights $w:V\rightarrow\mathbb{R}_{\geq 0}$. This is a special case of submodularity, so Theorem 3.1 gives a $(\nicefrac{{1}}{{2}}-\varepsilon)$-approximation. However, GIST achieves a stronger approximation ratio in this case.
Theorem 3.4.
Let $g(S)=\sum_{v\in S}w(v)$ be a linear function. For any $\varepsilon>0$, GIST outputs a set $S\subseteq V$ with $|S|\leq k$ and $f(S)\geq(\nicefrac{{2}}{{3}}-\varepsilon)\cdot\textnormal{OPT}$.
We defer the proof to Appendix B, but describe the main differences. For linear functions, Lemma 3.2 can be strengthened to show that GreedyIndependentSet outputs a set $T$ such that $g(T)\geq g(S_{d}^{*})$, for any $d^{\prime}<d/2$. Then, for the right distance threshold $d\in D$, we get two lower bounds $\textnormal{ALG}\geq\max\left\{g(S^{*})+d^{*}/(2(1+\varepsilon)),d^{*}\right\}$ that can be balanced appropriately.
Remark 3.5.
GIST iterates over $O(\log_{1+\varepsilon}(1/\varepsilon))$ distance thresholds. For each $d$, it inserts up to $k$ points with $O(n)$ distance and submodular valuation oracle calls per insertion, so these iterations consist of $O(nk\log_{1+\varepsilon}(1/\varepsilon))$ distance and submodular oracle calls. For linear functions, we can sort the points in advance by $w(v)$ and iterate over this ordering in all GreedyIndependentSet calls to reduce the number of valuation probes to $O(n\log_{1+\varepsilon}(1/\varepsilon))$.
4 Hardness of Approximation
[…]
Submodular utility.
There is no polynomial-time $0.5584$-approximation algorithm if $g$ is submodular.
Linear utility.
  • There is no polynomial-time $(\nicefrac{{2}}{{3}}+\varepsilon)$-approximation algorithm if $g$ is linear for general metrics, for any $\varepsilon>0$.
  • APX-completeness for linear utility functions even in the Euclidean metric, i.e., there is no polynomial-time approximation scheme (PTAS) for this problem.
4.1 General Metric Spaces
We build on the set cover hardness instances originally designed by Feige (1998); Feige et al. (2004); Feige and Vondrák (2010), and further engineered by Kapralov et al. (2013). In particular, we present the instances designed in Kapralov et al. (2013, Section 2) as follows:
Hardness of max $k$-cover.
For any constants $c,\varepsilon>0$ and a given collection/family of sets $\mathcal{F}$ partitioned into groups $\mathcal{F}_{1},\mathcal{F}_{2},\cdots,\mathcal{F}_{k}$, it is NP-hard to distinguish between the following two cases:
  • YES case: there exists $k$ disjoint sets $S_{i}$, one from each $\mathcal{F}_{i}$, whose union covers the entire universe.
  • NO case: for any $\ell\leq c\cdot k$ sets, their union covers at most a $(1-(1-\nicefrac{{1}}{{k}})^{\ell}+\varepsilon)$-fraction of the universe.
The construction above can be done such that each set $S_{i}\in\mathcal{F}=\bigcup_{r=1}^{k}\mathcal{F}_{r}$ has the same size. We also note that the parameter $k$ should be larger than any constant we set since there is always an exhaustive search algorithm with running time $|\mathcal{F}|^{k}$ to distinguish between the YES and NO cases. Therefore, we can assume, e.g., that $k\geq 1/\varepsilon$.
We use this max $k$-cover instance to prove hardness of approximation for the MDMS problem.
Theorem 4.1.
For any $\delta>0$, it is NP-hard to approximate MDMS within a factor of $\frac{2(1-1/e)}{2(1-1/e)+1}+\delta<0.55836+\delta$.
Proof.
We construct our instance based on the collection of sets above as follows. We set $\varepsilon=\min\{\delta,\delta^{2}/6\}$ and $c=1$. As mentioned above, we can assume that $k>\nicefrac{{1}}{{\varepsilon}}+1$ since there is always a brute force polynomial time algorithm for constant $k$ to distinguish between YES and NO instances.
Suppose there are $n$ sets $S_{1},S_{2},\cdots,S_{n}$ in the collection $\mathcal{F}$. We represent these $n$ sets with $n$ points in the MDMS instance. We overload the notation to show the corresponding point with $S_{i}$ too. For any subset of sets/points $T$, the submodular value $g(T)$ is defined as the cardinality of the union of corresponding sets, i.e., the coverage submodular function. In other words, $g(T)=|\bigcup_{S_{i}\in T}S_{i}|$.
The distances between points/sets are either $d$ or $2d$. If two sets are disjoint and belong to two separate partitions $\mathcal{F}_{r}$ and $\mathcal{F}_{r^{\prime}}$, their distance is $2d$. Otherwise, we set their distance to $d$. So for any two sets $S,S^{\prime}$ belonging to the same group $\mathcal{F}_{r}$, their distance is set to $d$. Also, for any two sets with nonempty intersection, we set their distance to $d$ too. Any other pair of points will have distance $2d$.
We set $d$ to be $(1-\nicefrac{{1}}{{e}})U$ where $U$ is the cardinality of the union of all sets, i.e., $U=|\bigcup_{S\in\mathcal{F}}S|$. In the YES case, the optimum solution is the family of $k$ disjoint sets from different groups that cover the entire universe. The value of optimum in this case is $2d+U=(2(1-\nicefrac{{1}}{{e}})+1)U$.
The algorithm has two possibilities: (case a) the algorithm gives up on the diversity objective and selects $k$ sets with minimum distance $d$, or (case b) it aims for a diversity term of $2d$. In case (a), the submodular value is at most $(1-(1-\nicefrac{{1}}{{k}})^{k}+\varepsilon)U$ because of the property of the NO case. We note that if the algorithm finds a set of $k$ sets with union size above this threshold, we can conclude that we have a YES instance, which contradicts the hardness result.
The limit of the upper bound for the submodular value as $k$ goes to infinity is $1-\nicefrac{{1}}{{e}}+\varepsilon$. We use its expansion series to derive:
$\left(1-\frac{1}{k}\right)^{k}\geq\frac{1}{e}-\frac{1}{2ek}-\frac{5}{24ek^{2}}-\cdots$
The sequence of negative terms declines in absolute value with a rate of at least $\nicefrac{{1}}{{k}}$. Thus, the absolute value of their total sum is at most the first deductive term $\nicefrac{{1}}{{2ek}}$ times $1/(1-\nicefrac{{1}}{{k}})=k/(k-1)$, which gives the simpler bound:
$\left(1-\frac{1}{k}\right)^{k}\geq\frac{1}{e}-\frac{1}{2e(k-1)}.$
Since $k-1$ is at least $\nicefrac{{1}}{{\varepsilon}}$, the submodular value in case (a) does not exceed:
$\left(1-\frac{1}{e}+\frac{1}{2e(k-1)}+\varepsilon\right)U\leq\left(1-\frac{1}{e}+2\varepsilon\right)U.$
Recall that $d=(1-\nicefrac{{1}}{{e}})U$, so the ratio of what the algorithm achieves in case (a) and the optimum solution of YES case is at most:
$\frac{(1-\nicefrac{{1}}{{e}}+2\varepsilon)U+d}{U+2d}$ $=\frac{2(1-\nicefrac{{1}}{{e}})+2\varepsilon}{2(1-\nicefrac{{1}}{{e}})+1}$
$\leq\frac{2(1-\nicefrac{{1}}{{e}})}{2(1-\nicefrac{{1}}{{e}})+1}+\delta.$
In case (b), the algorithm is forced to pick only disjoint sets to maintain a minimum distance of $2d$. This means if the algorithm picks $\ell$ sets, their union has size $\ell\cdot\nicefrac{{U}}{{k}}$. This is true because all sets in $\mathcal{F}$ have the same size and we know in the YES case, the union of $k$ disjoint sets covers the entire universe hence each set has size $\nicefrac{{U}}{{k}}$. For the special case of $\ell=1$, we know that only a $\nicefrac{{1}}{{k}}<\varepsilon$ fraction of universe is covered. We upper bound the covered fraction in terms of $\varepsilon$ for the other cases. The property of the NO case implies the following upper bound on $\ell$:
$\frac{\ell}{k}\leq 1-\left(1-\frac{1}{k}\right)^{\ell}+\varepsilon.$
We use the binomial expansion of $(1-\nicefrac{{1}}{{k}})^{\ell}$ and note that each negative term exceeds its following positive term in absolute value. This is true because of the cardinality constraint $\ell\leq k$. Therefore,
$\left(1-\frac{1}{k}\right)^{\ell}$ $\geq 1-\frac{\ell}{k}+\frac{\ell(\ell-1)}{2k^{2}}-\frac{\ell(\ell-1)(\ell-2)}{6k^{3}}$
$\geq 1-\frac{\ell}{k}+\frac{\ell(\ell-1)}{3k^{2}}$
$\geq 1-\frac{\ell}{k}+\frac{\ell^{2}}{6k^{2}},$
where the second to last inequality holds since $\ell-2<k$ and the last inequality holds because $\ell\geq 2$ and consequently $\ell-1\geq\nicefrac{{\ell}}{{2}}$. We can now revise the initial inequality:
$\frac{\ell}{k}$ $\leq(1-(1-\nicefrac{{1}}{{k}})^{\ell}+\varepsilon)$
$\leq 1-1+\frac{\ell}{k}-\frac{\ell^{2}}{6k^{2}}+\varepsilon\implies\frac{\ell^{2}}{6k^{2}}\leq\varepsilon.$
Thus, the fraction of the universe that the algorithm covers, namely $\nicefrac{{\ell}}{{k}}$, is at most $\sqrt{6\varepsilon}\leq\delta$.
In case (b), the ratio of what the algorithm achieves, and the optimum solution of YES case is at most:
$\frac{\delta\cdot U+2d}{U+2d}=\frac{2(1-\nicefrac{{1}}{{\mathrm{e}}})+\delta}{2(1-\nicefrac{{1}}{{\mathrm{e}}})+1}\leq\frac{2(1-\nicefrac{{1}}{{\mathrm{e}}})}{2(1-\nicefrac{{1}}{{\mathrm{e}}})+1}+\delta.$
This concludes the proof in both cases (a) and (b). ∎
4.2 Linear Utility
For the special case of linear utility functions, we prove a tight hardness-of-approximation result complementing the $\nicefrac{{2}}{{3}}$-approximation guarantee in Theorem 3.4.
Theorem 4.2.
[…]
We give the proof in Section C.1. Our construction builds on the work of Håstad (1996); Zuckerman (2006), which shows that the maximum clique problem does not admit an efficient $n^{1-\theta}$-approximation algorithm, for any constant $\theta>0$.
4.3 Euclidean Metric
[…]
Lemma 4.3 (Alimonti and Kann (2000, Theorem 3.2)).
The maximum independent set problem for graphs with degree at most $3$ is APX-complete.
Our second ingredient is a node embedding function $h_{G}(v)$ that encodes graph adjacency in Euclidean space.
Lemma 4.4.
Let $G=(V,E)$ be a simple undirected graph with $n=|V|$, $m=|E|$, and max degree $\Delta$. There exists an embedding $h_{G}:V\rightarrow\mathbb{R}^{n+m}$ such that if $\{u,v\}\in E$ then
$\lVert h_{G}(u)-h_{G}(v)\rVert_{2}\leq 1-\frac{1}{2(\Delta+1)},$
and if $\{u,v\}\not\in E$ then $\lVert h_{G}(u)-h_{G}(v)\rVert_{2}=1$.
Theorem 4.5.
The MDMS problem is APX-complete for the Euclidean metric even if $g$ is a linear function.
We sketch the main idea below and defer complete proofs to Appendix C. Let $\mathcal{I}(G)$ be the set of independent sets of $G$. Using Lemma 4.4, for any set of nodes $S\subseteq V$ with $|S|\geq 2$ in a graph $G$ with max degree $\Delta\leq 3$, we have the property:
  • $S\in\mathcal{I}(G)\implies\texttt{div}(S)=1$;
  • $S\not\in\mathcal{I}(G)\implies\texttt{div}(S)\leq 1-\frac{1}{2(\Delta+1)}\leq 1-\nicefrac{{1}}{{8}}$.
The gap between the cases is at least $\nicefrac{{1}}{{8}}$ for all such graphs (i.e., a universal constant), allowing us to prove that MDMS inherits the APX-completeness of bounded-degree max independent set.
5 Experiments
5.1 Synthetic Dataset
We use a collection of weighted random points and a budget-additive (submodular) utility function to (1) compare against baselines and (2) highlight the non-monotone behavior.
Setup.
Generate $n=1000$ points $\mathbf{x}_{i}\in\mathbb{R}^{d}$, for $d=64$, where each $\mathbf{x}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$ is i.i.d. and assigned a uniform continuous weight $w_{i}\sim\mathcal{U}_{[0,1]}$. For $\alpha,\beta\in[0,1]$, the objective function below trades off between the average (capped) utility of the selected points and their min-distance diversity:
$f(S)=\alpha\cdot\min\left\{\frac{1}{k}\sum_{i\in S}w_{i},\beta\right\}+(1-\alpha)\cdot\texttt{div}(S).$
We consider three baselines: random, simple, and greedy. random chooses a random set of $k$ points, permutes them, and returns the best prefix since the objective function is non-monotone. simple is the $0.387$-approximation algorithm in Section 2.1. greedy builds $S$ one point at a time by selecting the point with index $i^{*}_{t}=\operatorname*{arg\,max}_{i\in V\setminus S_{t-1}}f(S_{t-1}\cup\{i\})-f(S_{t-1})$ at each step $t\in[k]$, and outputs the best prefix $S=\operatorname*{arg\,max}_{t\in[k]}f(S_{t})$.
Results.
We plot the values of $f(S)$ for the baseline methods and GIST for all $k\in[n]$ in Figure 1. We set $\alpha=0.95$ to roughly balance the submodular and diversity terms. [^3] If $\mathbf{x},\mathbf{y}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$ are i.i.d., then $\mathbb{E}[\lVert\mathbf{x}-\mathbf{y}\rVert_{2}^{2}]=2d$. Setting $\frac{\alpha}{2}\approx(1-\alpha)\sqrt{2d}\implies\alpha\approx 1-\frac{1}{2\sqrt{2d}+1}=0.957$, for $\beta\geq 0.5$. We sweep over $\alpha$ and budget $\beta$ in Section D.1, and plot how these hyperparameters change $f(S)$. Note that the objective values are not directly comparable as a function of $k$ because of the $\nicefrac{{1}}{{k}}$ normalizing factor (hence the discontinuities), but both terms in the objective are decreasing on average in $k$ for a fixed $\alpha$. GIST dominates simple, greedy, and random in all cases. The poor performance of greedy is somewhat surprising since it is often competitive for subset selection tasks. simple beats greedy for all but the extreme values of $k$, i.e., small or close to $n$, but is always worse than GIST.
Figure 1: Baseline comparison plotting $f(S_{\text{ALG}})$ for $k\in[n]$ on synthetic data with $n=1000$, $\alpha=0.95$, and $\beta=0.75$.
5.2 Image Classification
Our ImageNet data sampling experiment compares the top-1 image classification accuracy achieved by different single-shot subset selection algorithms.
Setup.
We use the standard vision dataset ImageNet (Russakovsky et al., 2015) containing ~1.3 million images and 1000 classes. We select 10% of the points at random and use them to train an initial ResNet-56 model $\bm{\theta}_{0}$ (He et al., 2016). Then we use this initial model to compute an uncertainty value and 2048-dimensional embedding for each example. The margin-based uncertainty value of $\mathbf{x}_{i}$ is given by $u_{i}=1-(\Pr(y=b\mid\mathbf{x}_{i};~{}\bm{\theta}_{0})-\Pr(y=b^{\prime}\mid\mathbf{x}_{i};~{}\bm{\theta}_{0}))$, which measure the difference between the probability of the best and second-best class label $b^{\prime}$ for an example. Finally, we use the fast maximum inner product search of Guo et al. (2020) to build a $\Delta$-nearest neighbor graph in the embedding space using $\Delta=10$ and cosine distance. We present all model training hyperparameters in Appendix D.
We compare GIST with several state-of-the-art benchmarks:
  • random: We select samples from the dataset uniformly at random (without replacement). While this approach is simple and lightweight, it creates diversity in many settings and provides good solutions.
  • margin (Roth and Small, 2006): Margin sampling selects the top-$k$ points based on how hard they are to classify, i.e., using margin-based uncertainty values $u_{i}$.
  • $k$-center (Sener and Savarese, 2018): We run the classic greedy algorithm for $k$-center on the $\Delta$-nearest neighbor graph. The distance between non-adjacent nodes is implicitly assumed to be very large, which is reasonable in the large budget regime $k=\Omega(n)$.
  • submod: We select a subset by maximizing the submodular objective function $\alpha_{s}\sum_{i\in S}u_{i}-\beta_{s}\sum_{i,j\in S}s(i,j)$, under the budget constraint $|S|\leq k$. Similar pairwise submodular objective functions have been used in (Iyer et al., 2021; Kothawade et al., 2021; Ramalingam et al., 2021). Here $s(i,j)$ denotes the similarity between adjacent nodes using the same graph as in $k$-center. We tuned for best performance by selecting $\alpha_{s}=0.9$ and $\beta_{s}=0.1$.
We evaluate GIST using both linear and submodular utility functions. Let $f(S)=\alpha\cdot g(S)+(1-\alpha)\cdot\texttt{div}(S)$ for $\alpha\in[0,1]$. In Table 1, GIST-linear uses the same linear utility as margin: $g(S)=\sum_{i\in S}u_{i}$ with $\alpha=0.9$. GIST uses the same submodular utility as submod: $g(S)=\alpha_{s}\sum_{i\in S}u_{i}-\beta_{s}\sum_{i,j\in S}s(i,j)$ with $\alpha=0.95$, $\alpha_{s}=0.9$, and $\beta_{s}=0.1$. We use the same graph as $k$-center for distances.
Table 1: Top-1 classification accuracy (%) on ImageNet for different single-shot data downsampling algorithms. The cardinality constraint $k$ is expressed as a fraction of the ~1.3 million examples. The results are the average of 3 trials and the top performance for each $k$ is shown in bold.
$k$ (of data) random margin $k$-center submod GIST-linear GIST
30 66.27 65.97 66.70 66.52 66.89 67.23
40 69.17 69.73 70.06 70.71 70.51 70.75
50 71.05 72.33 73.09 72.69 72.67 73.14
60 72.49 73.42 73.60 74.20 74.37 74.30
70 73.70 74.49 74.24 75.24 75.41 75.32
80 74.42 75.01 75.17 75.91 75.96 75.45
90 75.16 75.10 74.97 75.84 76.12 76.03
Results.
We run each sampling algorithm with cardinality constraint $k$ on the full dataset to get a subset of examples that we then use to train a new ResNet-56 model. We report the top-1 classification accuracy of these models in Table 1. GIST with a linear or submodular utility is superior to all baselines. Interestingly, there is a cutover value of $k$ where GIST-linear becomes the best algorithm. Finally, observe that GIST-linear outperforms margin in all cases. This demonstrates how $\texttt{div}(S)$ encourages diversity in the subset of sampled points and ultimately improves downstream model quality.
Conclusion
This work presents a novel subset selection problem called MDMS that combines the total utility of the selected points (modeled by a monotone submodular function $g$) with the $\texttt{div}(S)=\min_{u,v\in S:u\neq v}\textnormal{dist}(u,v)$ diversity objective. We design and analyze the GIST algorithm, which achieves a $\nicefrac{{1}}{{2}}$-approximation guarantee by solving a series of maximum independent set instances on geometric intersection graphs with the GreedyIndependentSet bicriteria approximation algorithm. We complement our algorithm with a $0.5584$-hardness of approximation. From the theoretical perspective, it is an interesting open problem to bridge the gap between the algorithmic and hardness results of $0.5$ and $0.5584$ approximations.
For the special case of a linear utility function $g$, we show that GIST has a $\nicefrac{{2}}{{3}}$-approximation guarantee, and we complement this with a $(\nicefrac{{2}}{{3}}+\varepsilon)$-hardness of approximation, for any $\varepsilon>0$.
Finally, we present an empirical study showing that GIST steadily outperforms baseline methods for MDMS (in particular greedy) on a simple set of Gaussian points with random weights. We conclude by comparing the top-1 image classification accuracy achieved by GIST and three other state-of-the-art data sampling methods for ImageNet, demonstrating the benefit of optimizing for a blend of utility and diversity.
Appendix A $f(S)$ is Not Submodular
Consider the instance on four points $N=\{a,b,c,d\}\subseteq\mathbb{R}$ where $a=0$, $b=1$, and $c=d=2$ in the (one-dimensional) Euclidean metric, i.e., all the points are collinear. Let $g(S)=0$ for all $S\subseteq V$. For $f(S)$ to be submodular, it must hold that for every $S\subseteq T\subseteq N$ and $x\in N\setminus T$,
$f(S\cup\{x\})-f(S)\geq f(T\cup\{x\})-f(T).$
However, if $x=b$, $S=\{a,c\}$ and $T=\{a,c,d\}$, we have:
$f(S\cup\{x\})-f(S)$ $=f(\{a,b,c\})-f(\{a,c\})=1-2=-1$
$f(T\cup\{x\})-f(T)$ $=f(\{a,b,c,d\})-f(\{a,c,d\})=0-0=0,$
so $f(S)$ is not submodular.
In this instance, the function $f$ is not monotone. However, a similar monotone but still not submodular $f$ can be defined by setting $g(S)=\sum_{v\in S}w(v)$ where $w(v)=2$ for every $v\in N$.
Appendix B Proof of Theorem 3.4
Lemma B.1.
[…]
Proof.
Let $k^{\prime}=|T|$ and $t_{1},t_{2},\dots,t_{k^{\prime}}$ be the points in $T$ in the order that GreedyIndependentSet selected them. Let $B_{i}=\{v\in V:\textnormal{dist}(t_{i},v)<d^{\prime}\}$ be the points in $V$ contained in the radius-$d^{\prime}$ open ball around $t_{i}$. First, we show that each $B_{i}$ contains at most one point in $S_{d}^{*}$. If this is not true, then some $B_{i}$ contains two different points $u,v\in S_{d}^{*}$. Since $\textnormal{dist}(\cdot,\cdot)$ is a metric, this means
$\textnormal{dist}(u,v)$ $\leq\textnormal{dist}(u,t_{i})+\textnormal{dist}(t_{i},v)$
$<d^{\prime}+d^{\prime}$
$\leq d/2+d/2$
$=d,$
which contradicts the assumption that $S_{d}^{*}$ is an independent set of $G_{d}(V)$. Note that it is possible to have $B_{i}\cap B_{j}\neq\varnothing$, for $i\neq j$, since these balls consider all points in $V$.
Now let $C_{i}\subseteq V$ be the set of uncovered (by the open balls) points that become covered when GreedyIndependentSet selects $t_{i}$. Concretely, $C_{i}=B_{i}\setminus(B_{1}\cup\cdots\cup B_{i-1})$. Each $C_{i}$ contains at most one point in $S_{d}^{*}$ since $|B_{i}\cap S_{d}^{*}|\leq 1$. Moreover, if $s\in C_{i}\cap S_{d}^{*}$, then $w(t_{i})\geq w(s)$ because the points are sorted in non-increasing order and selected if uncovered.
Let $A=C_{1}\cup\cdots\cup C_{k^{\prime}}$ be the set of points covered by the algorithm. For each point $s\in S_{d}^{*}\cap A$, there is exactly one covering set $C_{i}$ corresponding to $s$. It follows that
$\sum_{s\in S_{d}^{*}\cap A}w(s)\leq\sum_{i\in[k^{\prime}]:S_{d}^{*}\cap C_{i}\neq\varnothing}w(t_{i}).$ (3)
It remains to account for the points in $S_{d}^{*}\setminus A$. Note that if we have any such points, then $|T|=k$ since the points in $S_{d}^{*}\setminus A$ are uncovered at the end of the algorithm. Further, for any $t_{i}\in T$ and $s\in S_{d}^{*}\setminus A$, we have $w(t_{i})\geq w(s)$ since $t_{i}$ was selected and the points are sorted by non-increasing weight. Therefore, we can assign each $s\in S_{d}^{*}\setminus A$ to a unique $C_{i}$ such that $C_{i}\cap S_{d}^{*}=\varnothing$. It follows that
$\sum_{s\in S_{d}^{*}\setminus A}w(s)\leq\sum_{i\in[k^{\prime}]:S_{d}^{*}\cap C_{i}=\varnothing}w(t_{i}).$ (4)
Adding the two sums together in Equation 3 and (4) completes the proof. ∎
See 3.4
Proof of Theorem 3.4.
Let $d^{*}$ be the minimum distance between two distinct points in $S^{*}$. There are two cases: $d^{*}<\varepsilon d_{\max}$ and $d^{*}\geq\varepsilon d_{\max}$. In the first case, outputting the $k$ heaviest points (Line 2) achieves a $(1-\varepsilon)$-approximation. To see this, first observe that
$\textnormal{OPT}\geq d_{\max}>\frac{d^{*}}{\varepsilon}\implies\varepsilon\cdot\textnormal{OPT}>d^{*}.$
The sum of the $k$ heaviest points upper bounds $g(S^{*})$, so we have
$\textnormal{ALG}\geq g(S^{*})=\textnormal{OPT}-d^{*}>(1-\varepsilon)\textnormal{OPT}.$
Therefore, we focus on the second case where $d^{*}\geq\varepsilon d_{\max}$. GIST considers a threshold $d=(1+\varepsilon)^{i}\cdot\varepsilon d_{\max}/2\in D$ such that
$\frac{d^{*}}{2}\in[d,(1+\varepsilon)d)\implies\frac{1}{1+\varepsilon}\cdot\frac{d^{*}}{2}<d\leq\frac{d^{*}}{2}.$
[…]
$f(T)$ $\geq g(S^{*})+d\geq g(S^{*})+\frac{d^{*}}{2(1+\varepsilon)}.$ (5)
The potential max-diameter update to $S$ on Line 5 gives us
$\textnormal{ALG}\geq d_{\max}\geq d^{*}.$ (6)
Putting (5) and (6) together, for any $0\leq p\leq 1$, it holds that
ALG $\geq p\cdot\left(g(S^{*})+\frac{d^{*}}{2(1+\varepsilon)}\right)+(1-p)\cdot d^{*}$
$=p\cdot g(S^{*})+\left(1-p+\frac{p}{2(1+\varepsilon)}\right)\cdot d^{*}.$
To maximize the approximation ratio as $\varepsilon\rightarrow 0$, we set $p=2/3$ by solving $p=1-p/2$. It follows that
ALG $\geq\frac{2}{3}\cdot g(S^{*})+\left(1-\frac{2}{3}+\frac{1}{3(1+\varepsilon)}\right)\cdot d^{*}$ (7)
$=\frac{2}{3}\cdot g(S^{*})+\frac{1}{3}\left(1+\frac{1}{1+\varepsilon}\right)\cdot d^{*}$
$\geq\frac{2}{3}\cdot g(S^{*})+\frac{1}{3}\left(2-\varepsilon\right)\cdot d^{*}$
$\geq\left(\nicefrac{{2}}{{3}}-\varepsilon\right)\cdot\textnormal{OPT},$
which completes the proof. ∎
Appendix C Missing Analysis for Section 4
C.1 Proof of Theorem: 4.2
See 4.2
Proof.
[…]
  • YES instance: graph $G$ has a clique of size $k$.
  • NO instance: graph $G$ does not have a clique of size greater than $\delta k$.
We reduce this instance of the maximum clique decision problem to MDMS with objective function (2) as follows. Represent each vertex of graph $G$ with a point in our ground set. The distance between a pair of points is $2$ if there is an edge between their corresponding vertices in $G$, and it is $1$ otherwise.
Use the same threshold value of $k$ (in the YES and NO instance above) for the cardinality constraint on set $S$, and set each weight $w(v)=1$. In a YES instance, selecting a clique of size $k$ as set $S$ results in the maximum possible value of the objective:
$\textnormal{OPT}=\alpha\cdot\frac{1}{k}\cdot k+(1-\alpha)\cdot 2=2-\alpha.$ (8)
In a NO instance, the best objective value that can be achieved in polynomial-time is the maximum of the following two scenarios: (a) selecting $k$ points with minimum distance $1$, or (b) selecting at most $\delta k$ vertices forming a clique with minimum distance $2$. The maximum value obtained by any polynomial-time algorithm is then
ALG $=\max\left\{\alpha+(1-\alpha)\cdot 1,\alpha\cdot\delta+(1-\alpha)\cdot 2\right\}$
$=\max\left\{1,2-(2-\delta)\alpha\right\}.$
These two terms become equal for $\alpha=1/(2-\delta)$. Thus, the gap between the maximum value any algorithm can achieve in the NO case and the optimum value in the YES case is
$\frac{1}{2-\alpha}=\frac{1}{2-\nicefrac{{1}}{{(2-\delta)}}}=\frac{2-\delta}{3-2\delta}.$
To complete the proof, it suffices to show that the ratio above is at most $\nicefrac{{2}}{{3}}+\varepsilon$. We separate the $\nicefrac{{2}}{{3}}$ term as follows:
$\frac{2-\delta}{3-2\delta}=\frac{\nicefrac{{2}}{{3}}\cdot(3-2\delta)+\nicefrac{{\delta}}{{3}}}{3-2\delta}=\frac{2}{3}+\frac{\delta}{9-6\delta}.$
Therefore, we must choose a value of $\delta$ satisfying $\nicefrac{{\delta}}{{(9-6\delta)}}\leq\varepsilon$. Since $\delta\leq 1$, the denominator $9-6\delta$ is positive. Equivalently, we want to satisfy:
$\frac{9-6\delta}{\delta}=\frac{9}{\delta}-6\geq\frac{1}{\varepsilon}.$
By setting $\delta<9\varepsilon/(1+6\varepsilon)$, we satisfy the required inequality and achieve the inapproximability gap in the theorem statement. ∎
C.2 Proof of Lemma 4.4
Let $G=(V,E)$ be a simple undirected graph. Our goal is to embed the vertices of $V$ into $\mathbb{R}^{d}$, for some $d\geq 1$, in a way that encodes the adjacency structure of $G$. Concretely, we want to construct a function $h_{G}:V\rightarrow\mathbb{R}^{d}$ such that:
  • $\lVert h_{G}(u)-h_{G}(v)\rVert_{2}=1$ if $\{u,v\}\not\in E$, and
  • $\lVert h_{G}(u)-h_{G}(v)\rVert_{2}\leq 1-\varepsilon_{G}$ if $\{u,v\}\in E$,
for the largest possible value of $\varepsilon_{G}\in(0,1]$.
Construction.
Let $n=|V|$ and $m=|E|$. Augment $G$ by adding a self-loop to each node to get $G^{\prime}=(V,E^{\prime})$. We embed $V$ using $G^{\prime}$ since each node now has positive degree. Let $\deg^{\prime}(v)$ be the degree of $v$ in $G^{\prime}$ and $N^{\prime}(v)$ be the neighborhood of $v$ in $G^{\prime}$.
Define a total ordering on $E^{\prime}$ (e.g., lexicographically by sorted endpoints $\{u,v\}$). Each edge $e\in E^{\prime}$ corresponds to an index in the embedding dimension $d\coloneqq|E^{\prime}|=m+n$. We consider the embedding function that acts as a degree-normalized adjacency vector:
$h_{G}(v)_{e}=\begin{cases}\sqrt{\frac{1}{2\deg^{\prime}(v)}}&\text{if $v\in e$},\\ 0&\text{if $v\not\in e$}.\end{cases}$ (9)
See 4.4
Proof.
If $\{u,v\}\not\in E$, then we have
$\lVert h_{G}(u)-h_{G}(v)\rVert_{2}^{2}$ $=\sum_{e\in N^{\prime}(u)}\left(\sqrt{\frac{1}{2\deg^{\prime}(u)}}-0\right)^{2}+\sum_{e\in N^{\prime}(v)}\left(\sqrt{\frac{1}{2\deg^{\prime}(v)}}-0\right)^{2}$
$=\left(\frac{1}{2}\sum_{e\in N^{\prime}(u)}\frac{1}{\deg^{\prime}(u)}\right)+\left(\frac{1}{2}\sum_{e\in N^{\prime}(v)}\frac{1}{\deg^{\prime}(v)}\right)$
$=\frac{1}{2}+\frac{1}{2}$
$=1.$
This follows because the only index where both embeddings can be nonzero is $\{u,v\}$, if it exists.
Now suppose that $\{u,v\}\in E$. It follows that
$\lVert h_{G}(u)-h_{G}(v)\rVert_{2}^{2}$ $=\sum_{e\in N^{\prime}(u)\setminus\{v\}}\frac{1}{2\deg^{\prime}(u)}+\sum_{e\in N^{\prime}(v)\setminus\{u\}}\frac{1}{2\deg^{\prime}(v)}+\left(\sqrt{\frac{1}{2\deg^{\prime}(u)}}-\sqrt{\frac{1}{2\deg^{\prime}(v)}}\right)^{2}$
$=\frac{\deg^{\prime}(u)-1}{2\deg^{\prime}(u)}+\frac{\deg^{\prime}(v)-1}{2\deg^{\prime}(v)}+\left(\frac{1}{2\deg^{\prime}(u)}+\frac{1}{2\deg^{\prime}(v)}-2\sqrt{\frac{1}{4\deg^{\prime}(u)\deg^{\prime}(v)}}\right)$
$=\frac{1}{2}+\frac{1}{2}-\sqrt{\frac{1}{\deg^{\prime}(u)\deg^{\prime}(v)}}$
$\leq 1-\frac{1}{\Delta+1}.$
The previous inequality follows from $\deg^{\prime}(v)=\deg(v)+1\leq\Delta+1$. For any $x\in[0,1]$, we have
$\sqrt{1-x}\leq 1-\frac{x}{2},$
so it follows that
$\lVert h_{G}(u)-h_{G}(v)\rVert_{2}\leq\sqrt{1-\frac{1}{\Delta+1}}\leq 1-\frac{1}{2(\Delta+1)},$
which completes the proof. ∎
C.3 Proof of Theorem 4.5
See 4.5
Proof.
[…]
  • YES instance: graph $G$ has an independent set of size $k$.
  • NO instance: graph $G$ does not have an independent set of size greater than $(1-\varepsilon_{0})k$.
We reduce this instance of bounded-degree maximum independent set to MDMS with objective function (2) as follows. Embed each node of the graph $G$ into Euclidean space using the function $h_{G}(v)$ in Lemma 4.4. We use the same threshold value of $k$ (between YES and NO instances above) for the cardinality constraint on set $S$, and we set each weight $w(v)=1$. In a YES instance, selecting an independent set of size $k$ as the set $S$ results in the maximum value of objective (2):
$\textnormal{OPT}=\alpha\cdot\frac{1}{k}\cdot k+(1-\alpha)\cdot 1=1,$
since $\lVert h_{G}(u)-h_{G}(v)\rVert_{2}=1$ for any two distinct points $u,v\in S$ since there is no edge between $u$ and $v$ in graph $G$.
In a NO instance, the best objective value that can be achieved in polynomial-time is the maximum of the following two scenarios: (a) selecting $k$ points with minimum distance at most $81-1/(2(\Delta+1))=1-1/$ 8, or (b) selecting at most $(1-\varepsilon_{0})k$ vertices forming an independent set with minimum distance equal to $1$. The maximum value obtained by any polynomial-time algorithm is then
ALG $=\max\left\{\alpha(1-\varepsilon_{0})+(1-\alpha)\cdot 1,\alpha+(1-\alpha)\left(1-1/8\right)\right\}$
$=\max\left\{1-\varepsilon_{0}\cdot\alpha,(7+\alpha)/8\right\}.$
These two terms become equal for $\alpha=1/(1+8\varepsilon_{0})$. Therefore, the gap between the maximum value any algorithm can achieve in the NO case and the optimum value in the YES case is upper bounded by
$1-\varepsilon_{0}\cdot\alpha=1-\frac{\varepsilon_{0}}{\nicefrac{{1}}{{\varepsilon_{0}}}+8}=1-\varepsilon_{1}.$
Since $\varepsilon_{0}>0$ is a constant, $\varepsilon_{1}\coloneqq\varepsilon_{0}/(\nicefrac{{1}}{{\varepsilon_{0}}}+8)>0$ is also a constant. This completes the proof of APX-completeness. ∎
Appendix D Additional Details for Section 5
D.1 Synthetic Dataset
We extend our comparison of baseline algorithms in Section 5.1 by sweeping over values of $\alpha,\beta$ in the objective function.
Figure 2: Baseline comparison with $\alpha\in(0.85,0.90,0.95,1.00)$ and $\beta=0.75$.
Figure 3: Baseline comparison with $\alpha=0.95$ and $\beta\in(0.60,0.70,0.80,0.90)$.
D.2 Image Classification
Hyperparameters for ImageNet classification.
We generate predictions and embeddings for all points using a coarsely-trained ResNet-56 model (He et al., 2016) trained on a random 10% subset of ImageNet (Russakovsky et al., 2015). We use SGD with Nesterov momentum 0.9 with 450/90 epochs. The base learning rate is 0.1, and is reduced by a tenth at 5, 30, 69, and 80. We extract the penultimate layer features to produce $2048$-dimensional embeddings of each image. We use the same hyperparameters as the original ResNet paper (He et al., 2016) with budgets and one-shot subset selection experiments designed in the same manner as Ramalingam et al. (2021).